# Mass, charge, and distance to Reissner-Nordström black hole in terms of directly measurable quantities

### v1

@@ Abstract
In this paper, we employ a general relativistic formalism and develop new theoretical tools that allow us to analytically express the mass and electric charge of the Reissner-Nordström black hole as well as its distance to a distant observer in terms of few directly observable quantities, such as the total frequency shift, aperture angle of the telescope, and redshift rapidity. Our analytic and concise formulas are valid on the midline, and the redshift rapidity is a relativistic invariant observable that represents the evolution of the frequency shift with respect to the proper time in the Reissner-Nordström spacetime. This procedure is applicable for particles undergoing circular motion around a spherically symmetric and electrically charged black hole, which is the case for accretion disks orbiting supermassive black holes hosted at the core of active galactic nuclei. Our results provide a novel method to measure the electric charge of the Reissner-Nordström black hole and its distance from the Earth, and the general formulas can be employed in black hole parameter estimation studies.
@@ I Introduction
Black holes are simple objects, completely described by only a handful of parameters, namely mass $M$, electric charge $Q$, and angular momentum $J$, known as the No-hair theorem nohair1 ; nohair3 ; nopelo . A black hole that contains all three parameters is called a Kerr-Newman black hole that reduces to a charged static Reissner-Nordström (hereafter RN) black hole in the spin-zero limit $J=0$. Although the electric charge in static RN black holes is often considered to be negligible RN , there is no conclusive evidence to assert that these bodies are not present in nature, hence electrically charged black holes have been extensively studied in the literature so far.
@@
It is worth mentioning that our analytic formulas enable us to obtain the RN black hole mass and charge as well as its distance from the Earth in a completely general relativistic framework. In this direction, one of our main goals in developing this approach to more general black hole solutions is improving the precision in measuring cosmic distances with respect to previous post-Newtonian methods that compute the angular-diameter distance to several megamaser systems, such as UGC 3789 MCP2 , NGC 6264 MCPV , NGC 5765b MCPVIII , NGC 4258 Humphreys2013 ; Reid2019 , CGCG 074-064 MCPXI .
The outline of the paper is as follows. In Sec. II, we elaborate on the geodesic motion of massive and null particles moving in the RN background. We derive expressions for the nonvanishing components of four-velocity of massive test particles and the photons’ four-momentum as well as for their light bending parameter in terms of RN black hole mass and charge. Next, in Sec. III, we present a brief overview of our general relativistic method in the RN background and use the results of Sec. II to obtain the frequency shifts of test particles circularly orbiting the RN black hole for an arbitrary point on the circular path. Then, we present a couple of expressions for the mass-to-distance ratio and charge-to-distance ratio of the RN black hole in terms of observational quantities on the midline. In Sec. IV, we define the redshift rapidity as the proper time evolution of the frequency shift in the RN background. Then, in Sec. V, we employ the results of the redshift and the redshift rapidity obtained, respectively, in Secs. III and IV, to express the black hole mass $M$, electric charge $Q$, and its distance $D$ from the Earth only in terms of observational elements. Finally, in Sec. VI, we conclude with some final remarks related to the developed general relativistic formalism as well as a brief discussion of the applications and relevance of our results.
@@ II Geodesic motion in Reissner-Nordström spacetime
In this section, we set the context of our general relativistic method and derive the nonvanishing components of the four-velocity of a test particle revolving the black hole spacetime in terms of the metric parameters by analyzing its equation of motion. Also, we obtain the components of the four-momentum for null particles and express the light bending parameter in terms of RN black hole mass and charge. The four-velocity elements of the massive test particles and their light bending parameter will be useful to construct an expression for the frequency shift in the RN background.
@@
The particles located in the vicinity of an RN black hole feel the curvature of spacetime produced by the black hole mass and charge. They keep the memory of the curvature and encode information of the RN black hole parameters in the frequency shift of emitted photons b ; AHerrera . Hence, further study regarding the motion of the massive/massless geodesic particles moving under the RN background (1) is necessary which we shall describe in the coming subsections.
@@ II.2 Geodesics of null particles
where $b_{\varphi}\equiv\frac{L_{\gamma}}{E_{\gamma}}$ is the bending of the light. By solving the previous relation for $b_{\varphi}$, we obtain an expression for the bending of the light experienced by photons emitted by test particles located at any point in a circular orbit. The light bending parameter in terms of the black hole parameters is given by
@@
It is worth mentioning that this parameter takes into account the bending of light due to the gravitational field in the vicinity of the RN black hole produced by mass and charge, and it is preserved along the whole null geodesics followed by photons from their emission till their detection. Therefore, since both $E_{\gamma}$ and $L_{\gamma}$ are constants of motion, we have $b_{\varphi,e}=b_{\varphi,d}$, where $b_{\varphi,e}$ refers to the light bending parameter when photons are emitted and $b_{\varphi,d}$ when they are detected.
@@ III Frequency shift in Reissner-Nordström background
Now, in order to be able to express the mass-to-distance ratio and the charge-to-distance ratio in terms of observables, let us take into account the values of the redshift and blueshift when the orbiting particles are placed on the midline where $\varphi=\pm\pi/2$, as shown in Fig. 3. At these points, the light bending parameter is maximized and Eq. (III) takes the following form
@@
$\tilde{Q}^{2}=\frac{2}{(R+B)^{2}}-\frac{3}{2}\Sigma+3\sec^{2}{\delta}-2,$ (50)
@@
for $\delta\ll 1$ on the midline. Then, we introduce the approximation (52) into Eqs. (49)-(50) and perform Taylor expansion about a small delta up to the third order to get
$\frac{M}{D}\approx\left(1-\frac{2}{RB}+\frac{4}{(R+B)^{2}}\right)\delta-\frac{(R-B)^{2}}{2RB}\delta^{3},$ (53)
$\frac{Q^{2}}{D^{2}}\approx\left(1-\frac{3}{RB}+\frac{8}{(R+B)^{2}}\right)\delta-\frac{3(R-B)^{2}}{4RB}\delta^{3},$ (54)
@@ IV Redshift rapidity in Reissner-Nordström background
since the temporal component $U^{t}_{e}$ (17) and azimuthal component $U^{\varphi}_{e}$ (18) of the four-velocity are constant quantities for circular orbits whereas the light bending parameter (36) depends on time through $\delta$ and $\varphi$. By making use of the chain rule, we have f
@@
where we used $U^{\varphi}_{e}=\frac{d\varphi}{d\tau}\Big{\rvert}_{r=r_{e}}$. Hence, by performing the derivatives of the light bending parameter (36) and the $\delta(\varphi)$-relation (39) presented in (58), the redshift rapidity for an arbitrary point on the circular orbit reads
@@ Acknowledgments
All authors are grateful to CONACYT for support under Grant No. CF-MG-2558591; M.M. also acknowledges SNI and was supported by CONACYT through the postdoctoral Grant No. 31155. A.H.-A. thanks SNI and PROMEP-SEP, and was supported by Grant VIEP-BUAP.

### v2

@@ Abstract
In this paper, we employ a general relativistic formalism and develop new theoretical tools that allow us to analytically express the mass and electric charge of the Reissner-Nordström black hole as well as its distance to a distant observer in terms of few directly observable quantities, such as the total frequency shift, aperture angle of the telescope, and redshift rapidity. Our analytic and concise formulas are valid on the midline, and the redshift rapidity is a relativistic invariant observable that represents the evolution of the frequency shift with respect to the proper time in the Reissner-Nordström spacetime. This procedure is applicable for particles undergoing circular motion around a spherically symmetric and electrically charged black hole, which is the case for accretion disks orbiting supermassive black holes hosted at the core of active galactic nuclei. Although this type of black hole is expected to be electrically neutral, our results provide a novel method to measure the electric charge of the Reissner-Nordström black hole, hence can shed some light on this claim. Besides, these results allow us to measure the mass of the black hole and its distance from the Earth, and we can employ the general formulas in black hole parameter estimation studies.
@@ I Introduction
Black holes are simple objects, completely described by only a handful of parameters, namely mass $M$, electric charge $Q$, and angular momentum $J$, known as the No-hair theorem nohair1 ; nohair3 ; nopelo . A black hole that contains all three parameters is called a Kerr-Newman black hole that reduces to a charged static Reissner-Nordström (hereafter RN) black hole in the spin-zero limit $J=0$.
Although the electric charge in static RN black holes is often considered to be negligible RN , there is no conclusive evidence to assert that these bodies are not present in nature (see ExtremeRN for observational signatures of the RN black hole), hence electrically charged black holes have been extensively studied in the literature so far . For instance, based on gravitational retro-lensing images, a method has been proposed to determine the electric charge of a RN black hole C1 , and to do so, the authors suggested employing the shadow size of the black hole. In addition, different methods for constraining the electric charge of the galactic center black hole Sgr A* are presented in C2 ; C3 ; C4 ; EHTimageConstraint . Despite presenting considerably good approximations for the black hole mass, there is no reliable value for the electric charge. However, it was claimed that a Reissner-Nordström black hole with a significant electric charge provides a better fit to the observational data for the black hole shadow in comparison with the Schwarzschild black hole C5 . In the present study, we propose a general relativistic method to determine the electric charge of an RN black hole and its mass in terms of directly measurable quantities.
@@
It is worth mentioning that our analytic formulas enable us to obtain the RN black hole mass and charge as well as its distance from the Earth in a completely general relativistic framework. In this direction, one of our main goals in developing this approach to more general black hole solutions is improving the precision in measuring cosmic distances with respect to previous post-Newtonian methods that compute the angular-diameter distance to several megamaser systems, such as UGC 3789 MCP2 , NGC 6264 MCPV , NGC 5765b MCPVIII , NGC 4258 Humphreys2013 ; Reid2019 , and CGCG 074-064 MCPXI .
The outline of the paper is as follows. In Sec. II, we elaborate on the geodesic motion of massive and null particles moving in the RN background. We derive expressions for the nonvanishing components of four-velocity of massive test particles and the photons’ four-momentum as well as for their impact parameter in terms of RN black hole mass and charge. Next, in Sec. III, we present a brief overview of our general relativistic method in the RN background and use the results of Sec. II to obtain the frequency shifts of test particles circularly orbiting the RN black hole for an arbitrary point on the circular path. Then, we present a couple of expressions for the mass-to-distance ratio and charge-to-distance ratio of the RN black hole in terms of observational quantities on the midline. In Sec. IV, we define the redshift rapidity as the proper time evolution of the frequency shift in the RN background. Then, in Sec. V, we employ the results of the redshift and the redshift rapidity obtained, respectively, in Secs. III and IV, to express the black hole mass $M$, electric charge $Q$, and its distance $D$ from the Earth only in terms of observational elements. Finally, in Sec. VI, we conclude with some final remarks related to the developed general relativistic formalism as well as a brief discussion of the applications and relevance of our results.
@@ II Geodesic motion in Reissner-Nordström spacetime
In this section, we set the context of our general relativistic method and derive the nonvanishing components of the four-velocity of a test particle revolving the black hole spacetime in terms of the metric parameters by analyzing its equation of motion. Also, we obtain the components of the four-momentum for null particles and express the impact parameter in terms of RN black hole mass and charge. The four-velocity elements of the massive test particles and their impact parameter will be useful to construct an expression for the frequency shift in the RN background.
@@
The particles located in the vicinity of an RN black hole feel the curvature of spacetime produced by the black hole mass and charge. They encode information about the spacetime curvature and the RN black hole parameters in the frequency shift of emitted photons b ; AHerrera . Hence, further study regarding the motion of the massive/massless geodesic particles moving under the RN background (1) is necessary which we shall describe in the coming subsections.
@@ II.2 Geodesics of null particles
where $b_{\varphi}\equiv\frac{L_{\gamma}}{E_{\gamma}}$ is the impact parameter RefResp . By solving the previous relation for $b_{\varphi}$, we obtain an expression for the bending of the light experienced by photons emitted by test particles located at any point in a circular orbit. The impact parameter in terms of the black hole parameters is given by
@@
It is worth mentioning that this parameter takes into account the bending of light due to the gravitational field in the vicinity of the RN black hole produced by mass and charge, and it is preserved along the whole null geodesics followed by photons from their emission till their detection. Therefore, since both $E_{\gamma}$ and $L_{\gamma}$ are constants of motion, we have $b_{\varphi,e}=b_{\varphi,d}$, where $b_{\varphi,e}$ refers to the impact parameter when photons are emitted and $b_{\varphi,d}$ when they are detected.
@@ III Frequency shift in Reissner-Nordström background
Now, in order to be able to express the mass-to-distance ratio and the charge-to-distance ratio in terms of observables, let us take into account the values of the redshift and blueshift when the orbiting particles are placed on the midline where $\varphi=\pm\pi/2$, as shown in Fig. 3. At these points, the impact parameter is maximized and Eq. (III) takes the following form
@@
$\tilde{Q}^{2}=\frac{8}{(R+B)^{2}}-\frac{3}{2}\Sigma+3\sec^{2}{\delta}-2,$ (50)
@@
for $\delta\ll 1$ on the midline. Then, we introduce the approximation (52) into Eqs. (49)-(50) and perform Taylor expansion about a small delta up to the fourth order to get
$\frac{M}{D}\approx$ $\left(1-\frac{2}{RB}+\frac{4}{(R+B)^{2}}\right)\delta$ (53)
$+\left(\frac{1}{3}-\frac{2}{3RB}+\frac{4}{3(R+B)^{2}}-\frac{(R-B)^{2}}{2RB}\right)\delta^{3},$
$\frac{Q^{2}}{D^{2}}\approx$ $\left(1-\frac{3}{RB}+\frac{8}{(R+B)^{2}}\right)\delta^{2}$ (54)
$+\left(\frac{2}{3}-\frac{2}{RB}+\frac{16}{3(R+B)^{2}}-\frac{3(R-B)^{2}}{4RB}\right)\delta^{4},$
@@ IV Redshift rapidity in Reissner-Nordström background
since the temporal component $U^{t}_{e}$ (17) and azimuthal component $U^{\varphi}_{e}$ (18) of the four-velocity are constant quantities for circular orbits whereas the impact parameter (36) depends on time through $\delta$ and $\varphi$. By making use of the chain rule, we have f
@@
where we used $U^{\varphi}_{e}=\frac{d\varphi}{d\tau}\Big{\rvert}_{r=r_{e}}$. Hence, by performing the derivatives of the impact parameter (36) and the $\delta(\varphi)$-relation (39) presented in (58), the redshift rapidity for an arbitrary point on the circular orbit reads
@@ VI Discussion and final remarks
As the final remark, we would like to mention that, in this study, our main goal was to obtain analytic formulas for the black hole mass and charge, and its distance from the Earth from a theoretical point of view. On the other hand, the general formulas of frequency shift and redshift rapidity can be used to estimate these parameters by employing observations from H2O megamaser systems on the accretion disks of supermassive black holes hosted at the core of AGNs in a similar manner as done for 17 supermassive Schwarzschild black holes towards ; Villalobos ; deby ; Villalobos2 . This is because, in such astrophysical systems, the emitter radius extends in the sub-parsec region from 0.04pc to 0.5pc from the center of the disk, hence located far from the black hole which allows this simple modeling to be applicable and robust. But in the case of the emitters located close enough to the black hole, one has to face additional difficulties, like the obfuscated determination of positions of photon sources due to the opacity size of the disk, and to incorporate the extreme conditions near the black hole, such as the consequences of the accretion process and the other properties of the disk, both in the theoretical description of the modeling as well as parameter estimation studies.
@@ Acknowledgments
All authors are grateful to CONACYT for support under Grant No. CF-MG-2558591; M.M. also acknowledges SNI and was supported by CONACYT through the postdoctoral Grants No. 31155 and No. 1242413. A.H.-A. thanks SNI and was supported by a VIEP-BUAP grant.